A large cylindrical tank contains 0.850 m3 of nitrogen gas at 28.0 ∘C and 8.50×103 Pa (absolute pressure). The tank has a tight-fitting piston that allows the volume to be change

What will be the pressure if the volume is decreased to 0.450m ^3 and the temperature is increased to 165 C?
Answer #1

from let P1,V1,T1 be the pressure ,volume and temperature in first state and  P2,V2,T2 be the pressure ,volume and temperature in second state

then from gas equation PV = nRT

P1*V1/T1 = P2*V2/T2

P2 = P1*(V1/V2)*(T2/T1)

= 8.50 x 10^3 x (0.850/0.450) x ((165 +273)/(28 +273)) Pa

=8.50 x 10^3 x 1.889 x 1.455 Pa

= 2.34 x 10^4 Pa